g(x)=x-2x^2
find the value of g(-3)+13

Answer by Timnewman(323)

hi dear,
If g(x)=x-2x^2
Then,
g(-3)+13=-3-2(-3)^2+13
=-3-2(9)+13
=-3-18+13
=-21+13
=-8
HOPE THIS HELPS?